Use the Well Ordering Principle to prove that any integer greater than or equal to 23 can be represented as the sum of nonnegative integer multiples of 6, 7 and 17.


1
Expert's answer
2021-11-02T18:35:17-0400

Let "S(n)" be the proposition that "n" can be represented as the sum of nonnegative integer multiples of 6, 7 and 17.

Let "C=\{n\geq 23\ |\ S(n) \ \text{is false}\ \}" .

We assume that "C" is nonempty.

Then by the Well Ordering Principle, there is a smallest number, "m\in C" .


"23=17+6\\ 24=6\cdot 4 \\ 25=6\cdot 3+7\\ 26=6\cdot 2+7\cdot 2
\\ 27=6+7\cdot 3
\\28=7\cdot 4"


Therefore, "m\geq29" .

Since "m>m-6\geq 23" , it follows that "(m-6)\not\in C" and "S(m-6)" is true.

But if "S(m-6)" is true, then "S(m)" must be true. Contradiction.

(if "n" can be represented as the sum of nonnegative integer multiples of 6, 7 and 17, then we can add 6 and get representation for number "n+6" )


So, "C" is empty and "S(n)" is true for "n\geq 23" .